# KeyDiff: Key Similarity-Based KV Cache Eviction for Long-Context LLM Inference in Resource-Constrained Environments

### v2

@@ B.1 Empirical Motivation for KeyDiff Setup
To construct Figure 5, we sample 100 prompts from the Qasper dataset in LongBench (Bai et al., 2024), compute the log determinant of $KK^{T}$ of the keys in the KV caches of each head and layer of Llama 3.2-3B-Instruct using a cache budget of $N=2048$ and a block size of $B=128$, and plot the distribution in Figure 5. We show this key distribution for sink attention, TOVA and KeyDiff.
B.2 Experiment Setup
In this subsection, we provide the experimental setup for KeyDiff and the baselines for the LongBench experiments. The LongBench evaluation is conducted using the default parameters of the LongBench evaluator with predefined prompt templates. The prompt block size is set to $B=128$, and the same $B$ value is used for all experiments.
For TOVA, H2O, ans SnapKV, the set of attention weights computed from a single key cache due to grouped query attention (Ainslie et al., 2023) is aggregated by taking the average over the attention weights. Additionally, only for SnapKV, we apply average smoothing to the attention score with a kernel size of 7 and keep the most recent 32 tokens in the cache, following the suggestion of the original paper. For Sink, we used the first four tokens as the attention sink, following the suggestion of the original paper.
B.3 Longbench dataset statistics
@@
We measure the number of tokens in the samples using LLama tokenizer (Touvron et al., 2023). As shown in Figure 10, LongBench exhibits variability in sample length from the datasets.
@@
The majority of other KV cache eviction studies assume an unconstrained memory footprint. Before they compress the cache by applying an eviction policy, they first set the target compression ratio and evict the appropriate number of KV pairs to satisfy the compression ratio (Zhang et al., 2024b; Oren et al., 2024; Li et al., 2024). On the other hand, we fix the KV cache size and ensure the number of cached tokens is less than or equal to the predefined cache size. Due to these differences, it is less straightforward to set appropriate KV cache budgets to satisfy the target compression ratios. Instead, we provide the average compression ratio, which is defined as:
@@ B.4 Additional Results
Single Doc. QA Multi Doc. QA Summarization Fewshot Learning Synthetic Code Narrative QA Qasper MF-en MF-zh HotpotQA 2WikiMQA Musique DuReader GovReport QMSum MultiNews VCSum TREC TriviaQA SAMSum LSHT PCount PR-en PR-zh Lcc RB-P Avg. Llama3.1-8B 30.05 47.00 56.12 59.86 57.33 47.81 32.25 35.64 34.86 25.32 27.02 17.28 73.00 91.61 43.37 45.50 8.33 99.50 99.00 61.66 51.94 49.74 H2O 2K 1.74 21.15 25.33 21.65 26.11 24.15 8.78 5.90 2.17 2.70 16.78 3.97 44.00 29.36 7.62 14.50 2.25 5.88 4.00 40.15 12.14 15.25 4K 4.07 36.16 36.00 38.02 33.52 32.87 17.78 5.68 6.66 5.95 24.09 6.03 55.00 47.65 17.41 18.50 4.00 24.50 31.25 54.85 21.43 24.83 6K 8.52 43.31 44.80 46.24 40.03 42.46 21.68 7.33 11.85 8.78 26.03 7.82 62.00 56.39 25.72 18.00 5.75 45.50 90.00 58.62 29.53 33.35 8K 13.85 44.94 47.81 56.14 43.64 44.90 23.65 11.01 18.78 11.35 26.49 9.96 69.50 69.05 33.41 19.50 5.25 62.50 98.67 59.74 36.26 38.40 TOVA 2K 22.57 37.26 39.43 36.96 45.74 34.48 14.77 16.98 28.87 21.17 26.95 16.21 62.50 90.73 42.74 18.75 0.00 18.00 32.00 62.68 52.48 34.35 4K 22.68 44.55 47.87 51.16 46.76 44.54 20.56 22.50 30.95 22.13 26.96 16.75 61.50 90.56 43.27 25.25 3.00 43.50 84.00 61.62 53.40 41.12 6K 24.59 45.93 53.92 55.45 55.09 47.43 25.07 27.68 32.33 24.10 27.00 16.91 68.50 90.81 43.89 29.00 4.25 67.00 98.67 61.50 52.39 45.31 8K 24.86 46.78 54.83 57.95 54.52 49.00 26.40 31.15 33.44 24.76 27.00 17.33 71.00 91.11 43.29 33.25 6.25 87.00 98.67 61.49 51.79 47.23 Sink 2K 21.83 34.27 29.24 32.82 38.64 29.50 12.59 16.18 28.51 20.21 26.62 15.54 65.00 89.46 42.20 22.25 2.00 25.50 32.50 64.95 59.54 33.78 4K 22.94 43.01 39.08 46.16 44.04 41.39 19.09 16.54 31.08 21.57 26.78 16.73 70.00 91.53 42.29 29.25 3.00 38.50 71.00 62.12 58.84 39.76 6K 25.41 47.40 44.13 52.78 47.39 45.73 21.90 17.55 32.53 22.19 26.87 17.05 72.00 91.25 43.41 33.75 3.08 52.50 98.00 62.22 56.24 43.49 8K 23.53 46.63 48.68 55.77 49.61 47.16 21.14 19.54 33.10 23.20 26.92 16.91 72.00 91.29 43.79 37.00 3.25 66.00 99.00 62.18 56.43 44.91 SnapKV 2K 21.81 37.22 37.19 38.29 46.10 35.42 16.53 16.37 29.83 21.05 26.77 16.16 61.00 88.84 42.56 21.75 4.03 51.50 81.17 62.37 51.45 38.45 4K 24.79 44.22 47.30 50.27 48.49 46.73 20.55 22.04 32.19 22.68 26.95 16.95 67.50 90.98 43.14 25.00 5.17 89.50 96.67 61.44 51.20 44.46 6K 24.10 45.57 50.44 55.27 53.12 48.41 24.27 27.46 33.43 23.53 27.03 16.84 71.50 92.28 43.58 27.00 5.25 98.00 99.00 61.32 52.16 46.65 8K 25.15 46.55 53.39 57.65 56.00 48.75 27.82 32.66 33.67 24.85 27.01 17.37 72.50 91.78 43.54 33.75 5.08 100.00 98.67 61.48 51.41 48.05 KeyL2Norm 2K 8.66 36.63 41.70 37.70 33.75 32.25 5.39 17.73 19.64 14.96 26.69 11.02 63.00 58.94 28.45 22.50 3.05 17.75 20.13 52.40 25.63 27.52 4K 15.38 44.06 51.75 47.52 50.22 45.56 18.44 27.61 29.50 22.27 26.93 13.44 69.50 79.41 37.50 27.00 4.50 58.00 80.50 58.82 35.08 40.14 6K 21.75 45.63 55.06 53.77 52.93 47.70 25.63 32.17 32.66 24.85 26.94 15.12 70.00 86.89 40.51 34.50 5.25 75.00 98.00 60.98 43.14 45.17 8K 25.12 45.70 56.02 56.57 58.14 47.77 30.29 34.09 33.81 24.89 26.94 15.89 71.50 89.26 41.34 39.00 7.25 87.00 99.00 62.05 48.28 47.61 KeyDiff 2K 26.64 41.73 50.99 51.18 51.59 46.47 22.84 32.37 29.02 23.86 26.76 14.81 66.50 85.92 39.26 42.25 3.17 96.00 96.25 59.17 39.42 45.06 4K 28.70 45.62 56.06 56.83 54.58 49.31 28.25 33.06 32.30 25.03 27.07 16.32 70.00 90.85 42.84 44.50 4.21 99.00 97.67 60.80 48.00 48.14 6K 29.90 46.33 55.11 59.00 56.80 49.50 31.52 34.97 33.44 24.58 26.98 16.80 72.00 90.99 43.10 47.00 5.27 99.50 99.00 61.40 49.70 49.19 8K 33.57 46.77 55.48 59.16 56.87 49.37 30.88 34.54 34.17 25.12 27.01 17.13 72.50 92.28 42.81 46.50 5.83 99.50 98.67 61.48 50.90 49.55 Llama3.2-3B 23.76 40.23 50.09 55.84 50.69 42.29 26.84 36.24 33.09 24.30 25.21 16.41 72.50 90.11 42.58 34.00 3.00 96.50 20.50 56.22 56.52 42.71 H2O 2K 1.63 19.96 20.20 15.20 18.02 19.56 2.88 6.47 0.78 1.55 15.97 3.11 41.00 21.97 9.83 11.75 0.50 0.50 0.00 39.71 13.91 12.60 4K 2.92 31.94 33.23 33.25 24.49 28.08 7.55 10.10 5.44 6.30 22.77 4.81 53.00 38.85 20.33 15.50 1.50 7.50 6.25 51.23 22.94 20.38 6K 4.62 38.81 39.06 45.17 34.66 35.52 15.21 13.36 10.51 10.01 24.25 6.66 61.50 53.23 27.37 15.25 0.50 13.00 19.50 54.55 32.29 26.43 8K 9.65 39.66 43.20 52.60 38.09 40.41 21.46 18.55 17.80 13.28 24.67 9.12 70.00 64.30 32.19 17.00 2.00 24.50 21.50 55.00 39.09 31.15 TOVA 2K 17.14 30.14 32.44 31.64 35.96 30.05 13.08 9.62 26.15 19.70 25.04 15.47 56.50 87.81 40.48 16.75 2.50 11.50 6.50 55.51 52.36 29.35 4K 20.52 39.53 42.47 45.80 44.12 38.42 18.22 17.76 29.36 21.36 24.96 16.60 63.50 88.98 41.50 18.75 3.00 23.50 15.00 55.72 56.66 34.56 6K 20.22 39.78 45.86 52.93 49.08 41.54 20.43 24.78 30.50 22.17 25.11 16.37 66.50 89.00 42.50 21.00 4.00 46.50 20.50 55.57 57.53 37.71 8K 21.08 40.67 49.07 55.17 48.69 41.93 23.05 31.02 31.64 22.85 25.21 16.55 69.00 89.25 42.19 22.50 2.50 71.00 21.50 55.77 57.47 39.91 Sink 2K 16.85 30.69 26.58 27.32 33.26 25.27 13.82 9.38 26.74 19.15 25.15 15.88 65.00 86.17 40.79 19.50 1.50 19.50 8.50 56.65 52.73 29.54 4K 19.46 38.61 36.22 41.68 41.97 35.84 13.37 9.86 29.34 20.19 25.06 16.44 71.00 88.06 41.31 21.75 2.50 35.50 16.00 56.48 52.43 33.96 6K 19.33 40.29 37.95 49.68 46.48 40.29 15.31 11.10 30.43 21.35 25.14 16.64 71.50 88.93 42.04 23.50 3.50 47.00 19.50 56.55 54.11 36.22 8K 20.15 40.02 41.94 53.57 48.15 42.24 16.01 14.76 31.64 22.10 25.20 16.50 73.00 89.26 42.37 26.25 3.50 62.50 20.50 56.86 56.63 38.25 SnapKV 2K 17.38 31.37 31.48 29.65 37.77 30.05 11.54 9.66 27.03 19.93 24.97 15.97 59.00 88.13 40.48 16.25 3.50 32.50 9.00 56.32 55.91 30.85 4K 19.85 39.22 39.86 47.33 46.70 37.98 16.64 16.88 29.79 21.21 25.01 16.74 65.50 89.35 40.95 18.25 2.50 62.50 22.50 55.74 56.88 36.73 6K 20.83 39.65 44.48 51.84 49.30 40.18 20.28 25.32 31.27 22.73 25.09 16.81 69.00 89.95 41.47 18.75 4.00 85.00 20.50 55.69 57.82 39.52 8K 20.49 40.80 48.16 55.44 48.78 41.65 24.79 30.40 31.81 23.46 25.17 16.44 70.00 90.17 41.99 22.00 5.00 94.00 21.50 55.77 57.29 41.20 KeyL2Norm 2K 7.67 30.39 31.85 30.64 29.47 25.76 7.41 14.17 15.36 12.42 24.20 7.91 48.00 50.99 23.09 17.50 2.00 7.50 5.00 48.92 26.32 22.22 4K 12.92 37.59 44.71 43.85 38.89 33.42 12.41 22.42 24.63 19.27 24.77 11.37 63.00 72.51 31.75 19.00 3.87 9.50 11.00 55.82 40.08 30.13 6K 13.02 40.55 48.17 50.87 43.18 40.17 17.10 31.29 28.99 21.47 25.08 13.30 65.00 79.61 37.16 21.50 2.50 46.50 18.50 56.49 47.05 35.60 8K 15.72 40.54 47.88 54.29 49.29 43.79 22.22 33.18 31.86 22.50 25.19 14.28 70.00 84.92 39.45 23.00 1.50 69.00 20.50 56.82 50.73 38.89 KeyDiff 2K 18.29 36.65 45.44 47.47 46.09 35.41 13.79 28.89 28.16 21.45 25.01 13.56 60.00 85.24 37.00 24.88 1.00 60.50 12.00 54.13 42.01 35.09 4K 22.34 40.60 49.15 52.56 50.14 40.30 21.65 32.46 31.38 23.44 25.06 15.28 66.50 87.92 41.41 27.50 2.50 88.50 19.50 55.55 52.24 40.28 6K 22.29 40.68 50.14 54.51 51.74 42.19 24.83 34.64 32.39 23.53 25.19 15.88 71.00 90.02 42.00 28.75 3.00 95.00 21.50 55.86 54.39 41.88 8K 22.41 40.77 50.10 55.62 49.83 43.58 28.09 34.30 32.78 23.60 25.17 15.77 72.00 90.17 42.46 30.75 3.50 96.50 21.50 55.85 55.65 42.40
@@
Appendix C Attention Sinks and Approximate Collinearity
(a) Keys
(b) Keys
(c) Queries
(d) Queries
Figure 11: Angular distributions of keys and queries measured with respect to to mean key and mean query
(a) Layer 25
(b) Layer 26
Figure 12: Cosine Similarity of mean key and mean query for each head in several layers.
To better understand why there exists a negative correlation between cosine similarity between keys and attention scores, we look to recent research that seeks to the importance of attention sinks in decoder-based LLMs. The authors in (Barbero et al., 2025) show that attention sinks emerge from training decoder-based LLMs since they can denoise the model and prevent rank collapse by limiting over mixing in attention heads. Moreover, attention patterns in decoder based models demonstrate that most attention logits are quite small (and almost always negative) for most keys and queries. This allows the attention sink to have high attention activation, preventing over mixing in addition to allowing the heads to specialize and selectively identify important tokens.
At the same time, the results in (Godey et al., 2024) suggest that hidden states, keys and queries are all approximately collinear in the sense that $\textrm{CosSim}(x_{i},x_{j})\gg 0$. In geometric terms, this means that most key tokens and query tokens lie in the same direction in Euclidean space. Our own results, seen in Figure 11 demonstrate that this is the case for both keys and queries. Our experiments show that most keys and queries lie within a small angular distance from the mean key and query. More than this, we see that across all heads, the mean key and mean query have negative cosine similarity Figure 12. Moreover, as seen in Figure 13, we find that the norms of keys and queries are tightly clustered around a relatively fixed value. This means that variations in the norm of key and query tokens have less impact on the magnitude of attention scores than their direction. These three observations indicate that most keys and queries combine to create uniformly small attention logits, and that larger attention weights are constructed by projecting keys closer to the direction of the mean query. This appears to be the fundamental mechanism through which over mixing is prevented: if most attention activations are very small, each head can increase the activations of a small number of keys across most queries selectively projecting them to be more aligned with the distribution of queries. This hypothesis is further supported by the fact that sink tokens themselves often have small norm, which results in an approximate no-op in the attention head as in (Barbero et al., 2025), however in this case, the only way for a key corresponding to a sink token to have high attention scores is if it is as parallel as possible to the set of query tokens.
(a) Keys
(b) Keys
(c) Queries
(d) Queries
Figure 13: Distribution of L2 norms for keys and queries in several layers.
Figure 14: Average cosine similarity per head between high scoring keydiff key tokens and mean query.
  • Most keys and queries have a similar L2 norm.
Theorem C.1.
Suppose that for a fixed query token $q$, there is a set of key tokens $\{k_{i}\}_{i=1}^{n}$ such that $||k_{i}||_{2}^{2}<M,\;\forall\;i$. Without loss of generality suppose $||q||=1$ and assume ${k^{*}}$ is a key not in $\{k_{i}\}_{i=1}^{n}$ with $||{k^{*}}||_{2}^{2}<M$ that has attention weight $w>0$:
Then
$\frac{\log(\frac{n}{n+1})-\log(1-w)}{2M}-1\leq\textrm{CosSim}({k^{*}},q)$
Proof.
To show this we have that
$w\sum\limits_{i=1}^{n}\exp(k_{i}^{\top}q)$ $=(1-w)\exp({k^{*}}^{\top}q)$
Note that $-M\leq k_{i}^{\top}q\leq M$ and hence
$w\sum\limits_{i=1}^{n}\exp(-M)$ $\leq(1-w)\exp({k^{*}}^{\top}q)$
$wn\exp(-M)$ $\leq(1-w)\exp({k^{*}}^{\top}q)$
$\frac{wn\exp(-M)}{1-w}$ $\leq\exp({k^{*}}^{\top}q)$
$\ \log(wn)-M-\log(1-w)$ $\leq M\textrm{CosSim}({k^{*}},q)$
$\frac{\log(wn)-M-\log(1-w)}{M}$ $\leq\textrm{CosSim}({k^{*}},q)$
$\frac{\log(\frac{n}{n+1})-2M-\log(1-w)}{M}$ $\leq 2\textrm{CosSim}({k^{*}},q)$
$\frac{\log(\frac{n}{n+1})-\log(1-w)}{2M}-1$ $\leq\textrm{CosSim}({k^{*}},q)$
In order to show this, we need the following auxiliary result.
Lemma C.2.
Proof.
$=\sum\limits_{i=1}^{n}\alpha_{i}x_{i}$
$=\sum\limits_{i=1}^{n}\langle\alpha_{i}x_{i},\alpha_{i}x_{i}\rangle$
$=\sum\limits_{i=1}^{n}\alpha_{i}^{2}$
proving the result. ∎
Theorem C.3.
Proof.
Consider the cosine similarity of $\bar{k}$ and ${k^{*}}$:
$\textrm{CosSim}(\bar{k},{k^{*}})$ $=\frac{{k^{*}}^{\top}\bar{k}}{||{k^{*}}||||\bar{k}||}$
$=\alpha_{q}\beta_{q}+\sum\limits_{i=1}^{n-1}\alpha_{i}\beta_{i}$
$\leq\alpha_{q}\beta_{q}+\sum\limits_{i=1}^{n-1}|\alpha_{i}||\beta_{i}|$
Applying Young’s inequality we obtain
$=\alpha_{q}\beta_{q}+\frac{1}{2}(1-\alpha_{q}^{2})+\frac{1}{2}(1-\beta_{q}^{2})$
$=1+\alpha_{q}\beta_{q}-\frac{1}{2}\alpha_{q}^{2}-\frac{1}{2}\beta_{q}^{2}$

### v3

@@ F.1 Empirical Motivation for KeyDiff Setup
To construct Figure 12, we sample 100 prompts from the Qasper dataset in LongBench [4], compute the log determinant of $KK^{T}$ of the keys in the KV caches of each head and layer of Llama 3.2-3B-Instruct using a cache budget of $N=2048$ and a block size of $B=128$, and plot the distribution in Figure 12. We show this key distribution for sink attention, TOVA and KeyDiff.
F.2 Experiment Setup
In this subsection, we provide the experimental setup for KeyDiff and the baselines for the LongBench experiments. The LongBench evaluation is conducted using the default parameters of the LongBench evaluator with predefined prompt templates. Tests are performed on NVIDIA A100 80GB GPUs.
For TOVA, H2O, and SnapKV, the set of attention weights computed from a single key cache due to grouped query attention [2] is aggregated by taking the average over the attention weights. Additionally, only for SnapKV, we apply average smoothing to the attention score with a kernel size of 7 and keep the most recent 32 tokens in the cache, following the suggestion of the original paper. For Sink, we used the first four tokens as the attention sink, following the suggestion of the original paper.
F.3 Longbench dataset statistics
@@
We measure the number of tokens in the samples using LLama tokenizer [31]. As shown in Figure 18, LongBench exhibits variability in sample length from the datasets.
@@
The majority of other KV cache eviction studies assume an unconstrained memory footprint. Before they compress the cache by applying an eviction policy, they first set the target compression ratio and evict the appropriate number of KV pairs to satisfy the compression ratio [43, 24, 20]. On the other hand, we fix the KV cache size and ensure the number of cached tokens is less than or equal to the predefined cache size. Due to these differences, it is less straightforward to set appropriate KV cache budgets to satisfy the target compression ratios. Instead, we provide the average compression ratio, which is defined as:
@@ F.4 Additional Results
Single Doc. QA Multi Doc. QA Summarization Fewshot Learning Synthetic Code Narrative QA Qasper MF-en MF-zh HotpotQA 2WikiMQA Musique DuReader GovReport QMSum MultiNews VCSum TREC TriviaQA SAMSum LSHT PCount PR-en PR-zh Lcc RB-P Avg. Llama3.1-8B 30.05 47.00 56.12 59.86 57.33 47.81 32.25 35.64 34.86 25.32 27.02 17.28 73.00 91.61 43.37 45.50 8.33 99.50 99.00 61.66 51.94 49.74 H2O 2K 1.74 21.15 25.33 21.65 26.11 24.15 8.78 5.90 2.17 2.70 16.78 3.97 44.00 29.36 7.62 14.50 2.25 5.88 4.00 40.15 12.14 15.25 4K 4.07 36.16 36.00 38.02 33.52 32.87 17.78 5.68 6.66 5.95 24.09 6.03 55.00 47.65 17.41 18.50 4.00 24.50 31.25 54.85 21.43 24.83 6K 8.52 43.31 44.80 46.24 40.03 42.46 21.68 7.33 11.85 8.78 26.03 7.82 62.00 56.39 25.72 18.00 5.75 45.50 90.00 58.62 29.53 33.35 8K 13.85 44.94 47.81 56.14 43.64 44.90 23.65 11.01 18.78 11.35 26.49 9.96 69.50 69.05 33.41 19.50 5.25 62.50 98.67 59.74 36.26 38.40 TOVA 2K 22.57 37.26 39.43 36.96 45.74 34.48 14.77 16.98 28.87 21.17 26.95 16.21 62.50 90.73 42.74 18.75 0.00 18.00 32.00 62.68 52.48 34.35 4K 22.68 44.55 47.87 51.16 46.76 44.54 20.56 22.50 30.95 22.13 26.96 16.75 61.50 90.56 43.27 25.25 3.00 43.50 84.00 61.62 53.40 41.12 6K 24.59 45.93 53.92 55.45 55.09 47.43 25.07 27.68 32.33 24.10 27.00 16.91 68.50 90.81 43.89 29.00 4.25 67.00 98.67 61.50 52.39 45.31 8K 24.86 46.78 54.83 57.95 54.52 49.00 26.40 31.15 33.44 24.76 27.00 17.33 71.00 91.11 43.29 33.25 6.25 87.00 98.67 61.49 51.79 47.23 Sink 2K 21.83 34.27 29.24 32.82 38.64 29.50 12.59 16.18 28.51 20.21 26.62 15.54 65.00 89.46 42.20 22.25 2.00 25.50 32.50 64.95 59.54 33.78 4K 22.94 43.01 39.08 46.16 44.04 41.39 19.09 16.54 31.08 21.57 26.78 16.73 70.00 91.53 42.29 29.25 3.00 38.50 71.00 62.12 58.84 39.76 6K 25.41 47.40 44.13 52.78 47.39 45.73 21.90 17.55 32.53 22.19 26.87 17.05 72.00 91.25 43.41 33.75 3.08 52.50 98.00 62.22 56.24 43.49 8K 23.53 46.63 48.68 55.77 49.61 47.16 21.14 19.54 33.10 23.20 26.92 16.91 72.00 91.29 43.79 37.00 3.25 66.00 99.00 62.18 56.43 44.91 SnapKV 2K 21.81 37.22 37.19 38.29 46.10 35.42 16.53 16.37 29.83 21.05 26.77 16.16 61.00 88.84 42.56 21.75 4.03 51.50 81.17 62.37 51.45 38.45 4K 24.79 44.22 47.30 50.27 48.49 46.73 20.55 22.04 32.19 22.68 26.95 16.95 67.50 90.98 43.14 25.00 5.17 89.50 96.67 61.44 51.20 44.46 6K 24.10 45.57 50.44 55.27 53.12 48.41 24.27 27.46 33.43 23.53 27.03 16.84 71.50 92.28 43.58 27.00 5.25 98.00 99.00 61.32 52.16 46.65 8K 25.15 46.55 53.39 57.65 56.00 48.75 27.82 32.66 33.67 24.85 27.01 17.37 72.50 91.78 43.54 33.75 5.08 100.00 98.67 61.48 51.41 48.05 KeyL2Norm[9] 2K 8.66 36.63 41.70 37.70 33.75 32.25 5.39 17.73 19.64 14.96 26.69 11.02 63.00 58.94 28.45 22.50 3.05 17.75 20.13 52.40 25.63 27.52 4K 15.38 44.06 51.75 47.52 50.22 45.56 18.44 27.61 29.50 22.27 26.93 13.44 69.50 79.41 37.50 27.00 4.50 58.00 80.50 58.82 35.08 40.14 6K 21.75 45.63 55.06 53.77 52.93 47.70 25.63 32.17 32.66 24.85 26.94 15.12 70.00 86.89 40.51 34.50 5.25 75.00 98.00 60.98 43.14 45.17 8K 25.12 45.70 56.02 56.57 58.14 47.77 30.29 34.09 33.81 24.89 26.94 15.89 71.50 89.26 41.34 39.00 7.25 87.00 99.00 62.05 48.28 47.61 KeyDiff 2K 26.64 41.73 50.99 51.18 51.59 46.47 22.84 32.37 29.02 23.86 26.76 14.81 66.50 85.92 39.26 42.25 3.17 96.00 96.25 59.17 39.42 45.06 4K 28.70 45.62 56.06 56.83 54.58 49.31 28.25 33.06 32.30 25.03 27.07 16.32 70.00 90.85 42.84 44.50 4.21 99.00 97.67 60.80 48.00 48.14 6K 29.90 46.33 55.11 59.00 56.80 49.50 31.52 34.97 33.44 24.58 26.98 16.80 72.00 90.99 43.10 47.00 5.27 99.50 99.00 61.40 49.70 49.19 8K 33.57 46.77 55.48 59.16 56.87 49.37 30.88 34.54 34.17 25.12 27.01 17.13 72.50 92.28 42.81 46.50 5.83 99.50 98.67 61.48 50.90 49.55 Llama3.2-3B 23.76 40.23 50.09 55.84 50.69 42.29 26.84 36.24 33.09 24.30 25.21 16.41 72.50 90.11 42.58 34.00 3.00 96.50 20.50 56.22 56.52 42.71 H2O 2K 1.63 19.96 20.20 15.20 18.02 19.56 2.88 6.47 0.78 1.55 15.97 3.11 41.00 21.97 9.83 11.75 0.50 0.50 0.00 39.71 13.91 12.60 4K 2.92 31.94 33.23 33.25 24.49 28.08 7.55 10.10 5.44 6.30 22.77 4.81 53.00 38.85 20.33 15.50 1.50 7.50 6.25 51.23 22.94 20.38 6K 4.62 38.81 39.06 45.17 34.66 35.52 15.21 13.36 10.51 10.01 24.25 6.66 61.50 53.23 27.37 15.25 0.50 13.00 19.50 54.55 32.29 26.43 8K 9.65 39.66 43.20 52.60 38.09 40.41 21.46 18.55 17.80 13.28 24.67 9.12 70.00 64.30 32.19 17.00 2.00 24.50 21.50 55.00 39.09 31.15 TOVA 2K 17.14 30.14 32.44 31.64 35.96 30.05 13.08 9.62 26.15 19.70 25.04 15.47 56.50 87.81 40.48 16.75 2.50 11.50 6.50 55.51 52.36 29.35 4K 20.52 39.53 42.47 45.80 44.12 38.42 18.22 17.76 29.36 21.36 24.96 16.60 63.50 88.98 41.50 18.75 3.00 23.50 15.00 55.72 56.66 34.56 6K 20.22 39.78 45.86 52.93 49.08 41.54 20.43 24.78 30.50 22.17 25.11 16.37 66.50 89.00 42.50 21.00 4.00 46.50 20.50 55.57 57.53 37.71 8K 21.08 40.67 49.07 55.17 48.69 41.93 23.05 31.02 31.64 22.85 25.21 16.55 69.00 89.25 42.19 22.50 2.50 71.00 21.50 55.77 57.47 39.91 Sink 2K 16.85 30.69 26.58 27.32 33.26 25.27 13.82 9.38 26.74 19.15 25.15 15.88 65.00 86.17 40.79 19.50 1.50 19.50 8.50 56.65 52.73 29.54 4K 19.46 38.61 36.22 41.68 41.97 35.84 13.37 9.86 29.34 20.19 25.06 16.44 71.00 88.06 41.31 21.75 2.50 35.50 16.00 56.48 52.43 33.96 6K 19.33 40.29 37.95 49.68 46.48 40.29 15.31 11.10 30.43 21.35 25.14 16.64 71.50 88.93 42.04 23.50 3.50 47.00 19.50 56.55 54.11 36.22 8K 20.15 40.02 41.94 53.57 48.15 42.24 16.01 14.76 31.64 22.10 25.20 16.50 73.00 89.26 42.37 26.25 3.50 62.50 20.50 56.86 56.63 38.25 SnapKV 2K 17.38 31.37 31.48 29.65 37.77 30.05 11.54 9.66 27.03 19.93 24.97 15.97 59.00 88.13 40.48 16.25 3.50 32.50 9.00 56.32 55.91 30.85 4K 19.85 39.22 39.86 47.33 46.70 37.98 16.64 16.88 29.79 21.21 25.01 16.74 65.50 89.35 40.95 18.25 2.50 62.50 22.50 55.74 56.88 36.73 6K 20.83 39.65 44.48 51.84 49.30 40.18 20.28 25.32 31.27 22.73 25.09 16.81 69.00 89.95 41.47 18.75 4.00 85.00 20.50 55.69 57.82 39.52 8K 20.49 40.80 48.16 55.44 48.78 41.65 24.79 30.40 31.81 23.46 25.17 16.44 70.00 90.17 41.99 22.00 5.00 94.00 21.50 55.77 57.29 41.20 KeyL2Norm [9] 2K 7.67 30.39 31.85 30.64 29.47 25.76 7.41 14.17 15.36 12.42 24.20 7.91 48.00 50.99 23.09 17.50 2.00 7.50 5.00 48.92 26.32 22.22 4K 12.92 37.59 44.71 43.85 38.89 33.42 12.41 22.42 24.63 19.27 24.77 11.37 63.00 72.51 31.75 19.00 3.87 9.50 11.00 55.82 40.08 30.13 6K 13.02 40.55 48.17 50.87 43.18 40.17 17.10 31.29 28.99 21.47 25.08 13.30 65.00 79.61 37.16 21.50 2.50 46.50 18.50 56.49 47.05 35.60 8K 15.72 40.54 47.88 54.29 49.29 43.79 22.22 33.18 31.86 22.50 25.19 14.28 70.00 84.92 39.45 23.00 1.50 69.00 20.50 56.82 50.73 38.89 KeyDiff 2K 18.29 36.65 45.44 47.47 46.09 35.41 13.79 28.89 28.16 21.45 25.01 13.56 60.00 85.24 37.00 24.88 1.00 60.50 12.00 54.13 42.01 35.09 4K 22.34 40.60 49.15 52.56 50.14 40.30 21.65 32.46 31.38 23.44 25.06 15.28 66.50 87.92 41.41 27.50 2.50 88.50 19.50 55.55 52.24 40.28 6K 22.29 40.68 50.14 54.51 51.74 42.19 24.83 34.64 32.39 23.53 25.19 15.88 71.00 90.02 42.00 28.75 3.00 95.00 21.50 55.86 54.39 41.88 8K 22.41 40.77 50.10 55.62 49.83 43.58 28.09 34.30 32.78 23.60 25.17 15.77 72.00 90.17 42.46 30.75 3.50 96.50 21.50 55.85 55.65 42.40
@@
Appendix G Ablation study
Table 14: Anchor vector ablation study. Average of Full Longbench results for Llama 3.2-3B-Instruct. KeyDiff results match Table 10. (Higher is better)
2K 4K 6K 8K
KeyDiff 35.09 40.28 41.88 42.40
$\operatorname{Pairwise}$ 35.24 40.61 41.87 42.45
$\operatorname{Median}$ 35.43 40.67 41.89 42.26
Selecting the Anchor Vector
Table 15: Distance metric ablation study. Average of Full Longbench results of Llama 3.2-3B-Instruct. KeyDiff results match Table 10. (Higher is better)
2K 4K 6K 8K
KeyDiff 35.09 40.28 41.88 42.40
$\operatorname{DotProd}$ 30.14 38.01 41.09 42.23
$\operatorname{Euclidean}$ 13.68 21.06 25.91 29.53
Selecting the Similarity Metric
G.1 Needle in a Haystack results for Sink attention
Figure 19: Needle in a Haystack results for Sink Attention [35]